OpenStudy (anonymous):

lim (x→3) (5^4x−5^12 )/ (3−x) is equal to the derivative of a function f at a point a (that is, f′(a)). What is the function f in this case?

OpenStudy (anonymous):

it could most probably be \(5^{4x}\) but in that case, the denominator would be x-3 and not 3-x

OpenStudy (anonymous):

yeah that's not the answer but thx :P

OpenStudy (amistre64):

so its giving you a derivative (the long way) and asking what function its deriving

OpenStudy (anonymous):

yes but i don't know what the original function is

OpenStudy (amistre64):

\[\frac{f(x)-f(a)}{x-a}\] \[\frac{5^{4x}-5^{4a}}{x-a}\] \[\frac{5^{4x}-5^{12}}{x-3}\] \[-\frac{5^{4x}-5^{12}}{3-x}\] so to keep the parts in place, we are missing a negative arent we

OpenStudy (anonymous):

but what is the original function? i got -5^4x but it's only half right!

OpenStudy (anonymous):

okay never mind i got it!

OpenStudy (amistre64):

thats the original function, whats the other half of the question?

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

youre welcome